# FFHFlow: A Flow-based Variational Approach for Learning Diverse Dexterous Grasps with Shape-Aware Introspection

### v2

@@ IV-C1 Learning Grasp Distribution via DLVMs
With Jensen Inequality, we have the following variational lower bound with an approximate posterior of the latent variable $q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{g})$ (more derivation details in appendix):
$\begin{split}\log{p_{\theta}(\mathbf{g}|\mathbf{x})}\geq&\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{g})}[\log{p_{\theta}(\mathbf{g}|\mathbf{x},\mathbf{z})}]-\\ &\beta KL(q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{g})||p_{\theta}(\mathbf{z}|\mathbf{x})).\end{split}$ (3)
@@
where $\beta$ is a hyperparameter. In practice, the KLD term is implemented as the sum of the entropy of the approximate posterior and the cross-entropy between the prior and the approximate posterior for convenience. The effect of minimizing this term is to match the input-dependent prior $p_{\theta}(\mathbf{z}|\mathbf{x})$ to the approximate posterior $q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{g})$, and meanwhile encourage the approximate posterior to be more diverse. A flexible prior is of vital importance at inference time as we need to resort to the prior for drawing latent samples. These samples will be conditioned on the Grasp Flow for grasp generation (see Figure 2). Therefore, an input-dependent and learnable prior is not only useful for expressivity increase but also necessary for generating diverse grasps given an incomplete point cloud.
Grasp Flow $p_{\theta}(\mathbf{g}|\mathbf{x})$
@@
Both VAE and GANs are reported to suffer from the mode collapse problem [5], which is also named Information Preference Property in the literature [4]. It illustrates the inability of the VAE to capture the entire data distribution, in our case, the complex multi-modal grasp distribution. Whenever this happens, the latent variable $\mathbf{z}$ is neglected by the powerful decoder and hence is uninformative in terms of the data, i.e., , grasps and object point clouds. It has been proved that the unbounded likelihood function is the crux of this problem, which is usually an isotropic Gaussian observation model [3].
To mitigate this, we propose to use cNF for the likelihood function $p_{\theta}(\mathbf{g}|\mathbf{x},\mathbf{z})$ by abuse of notations in Section IV-B: $p_{\theta}(\mathbf{g}|\mathbf{x},\mathbf{z})=p_{u|\mathbf{z}}(T_{\theta|\mathbf{z}}^{-1}(\mathbf{g};\mathbf{z});\mathbf{z})|\det(J_{T_{\theta|\mathbf{z}}^{-1}}(\mathbf{g};\mathbf{z}))|$, where the base distribution $p_{u|\mathbf{z}}:\mathcal{R}^{d}\times\mathcal{R}^{l}\rightarrow\mathcal{R}$ is conditional. As the latent variable $\mathbf{z}$ subsumes the information from $\mathbf{x}$, $\mathbf{x}$ is not shown on the RHS for conciseness. In contrast to cVAE, NFs allow the model to learn a more general likelihood function instead of an isotropic Gaussian as there is no underlying assumption for the distribution form. Meanwhile, the architecture of NFs is restricted to be a diffeomorphism, which may help alleviate the unboundness issue.
Prior Flow $p_{\theta}(\mathbf{z}|\mathbf{x})$
@@
The overly-simple prior can induce excessive regularization, limiting the quality of the latent representation [6, 7], which can be observed in cVAE-based approaches [1, 12, 2] with an input-independent isotropic Gaussian as the prior (Figure 3b). On the other hand, the assumption of input-independence for the prior [4] poses difficulty in learning informative latent features.
To address this, we propose to utilize a second cNF for an input-dependent prior distribution, in our case, a point cloud-dependent prior: $p_{\theta}(\mathbf{z}|\mathbf{x})=p_{u}(T_{\theta|\mathbf{x}}^{-1}(\mathbf{z};\mathbf{x}))|\det(J_{T_{\theta|\mathbf{x}}^{-1}}(\mathbf{z};\mathbf{x}))|$.
Variational Inference Network $q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{g})$
[…]
Grasp Synthesis and Shape-Aware Introspection
@@
Besides improving the quality of the latent features, an object point-cloud-dependent prior offers a novel way to grasp synthesis at inference time. For a test object point cloud $\mathbf{x}^{*}$, we can generate the corresponding grasps $\mathbf{g}^{*}$ by performing ancestral sampling:
@@
From section III we know that we are able to compute the exact likelihoods of $p_{\theta}(\mathbf{g}|\mathbf{z}^{*})$ and $p_{\theta}(\mathbf{z}|\mathbf{x}^{*})$ from Grasp Flow and Prior Flow, respectively. We expect these two quantities to capture the knowledge gained by the model quantitatively. With these, our model can be introspective against its unknown knowledge, such as the incomplete shapes due to partial observation and unknown object shapes due to limited coverage of training data. Specifically, the likelihoods of Grasp Flow quantify the view-level shape uncertainty as this flow considers both the point cloud and the corresponding grasps. Complementarily, thanks to the expressive latent variables, Prior Flow can represent the object-level shape uncertainty.
IV-D Uncertainty-Aware Grasp Evaluation
@@
To further secure the grasp success, we train a discriminative grasp evaluator $f_{\psi}(\mathbf{g},\mathbf{x})$ that outputs a score to better capture the grasp quality by learning to distinguish feasible grasps from infeasible ones in a supervised manner [12]. Furthermore, as the Grasp Flow is able to quantify the uncertainty of incomplete shapes, it is beneficial to incorporate this information into the grasp evaluation. With this, we can penalize the grasps approaching the uncertain side of the partially observed object. To this end, we propose an uncertainty-aware grasp evaluation scheme that fuses the batch-normalized view-level shape uncertainty into the grasp evaluator:
$\epsilon f_{\psi}(\mathbf{g},\mathbf{x})+(1-\epsilon)\log{p_{\theta}(\mathbf{g}|\mathbf{z}^{*})}$ (5)
, where $\epsilon\in[0,1]$ is a parameter for balancing the effects of increasing grasp quality and lowering uncertainty.
V Experiment
@@
In this section, we perform comprehensive experiments and analysis to answer the following questions: Q1: Can the proposed FFHFlow-lvm facilitate more expressive latent representations for both diverse, high-quality and fast grasp synthesis? Q2: How well does FFHFlow-lvm perform compared with other state-of-the-art grasp synthesis approaches? Q3: How effective is the shape-aware introspection of FFHFlow-lvm against OOD data and does it benefit grasp evaluation? Q4: Can the model generalize to the real world for single object grasping in free and confined space, as well as cluttered scenes?
TABLE I: Average Success Rate and Run-time in Simulation
Methods Similar Novel Run time (ms) Heuristic $20.9\%$ $11.1\%$ $3387$ cVAE [12] $84.6\%$ $52.4\%$ 30 GAN [20] $86.0\%$ $49.4\%$ 30 Diffusion [8] $88.2\%$ $51.7\%$ $1610$ FFHFlow-lvm 94.6% 52.7% $130$
Figure 4: Objects and Setups Evaluated in the Simulation and Real World. We test simulation objects with similar (a) and novel (b) shapes, real objects (c), and grasping in confined space (d) and clutter (e-f).
V-A Experimental Setup
@@
The experiments of grasping single tabletop objects are performed in both the simulation and the real world with the DLR-HIT II hand [53]. In the real-world experiments, we validate our approach in two more complex ones, i.e., the confined and cluttered scenarios. The success criterion is defined as the ability to lift the object 20 cm above its resting position without slippage.
V-B Implementation Details
Based on the normflows package [54], we implement FFHFlow-lvm with an 8-layered conditional Glow [36] where each layer has a 4-layered Multi-Layer Perceptron (MLP) for predicting the parameters of the affine operation. The model is trained with a learning rate of $1\mathrm{e}{-4}$ and a mini-batch size of $64$ for $16$ epochs or $20k$ iterations. The objective in Equation 3 is then optimized with a linearly increased $\beta$ from $1\mathrm{e}{-7}$ to $1\mathrm{e}{-1}$ in each iteration with AdamW optimizer. Inspired by [55], to capture the sensitive high-frequency contents in grasping poses, we additionally use positional encoding for grasp poses.
V-C Datasets
We use only simulated data generated based on a heuristic grasp planner for training in Gazebo Simulator [12]. For training, we use 75 graspable objects filtered from KIT [56] datasets based on their graspability and object type. For testing, we select objects of two levels of difficulty from the KIT and YCB dataset [57] in simulation (Figure 4):
  • Similar: 11 objects from KIT dataset with geometries similar to those in the training set.
  • Novel: 9 objects from YCB dataset with geometries distinct from those in the training set, often more difficult to grasp.
For real-world evaluation, we use 11 unknown objects from YCB dataset. Additionally, we evaluate more complex real-world settings, i.e., confined and cluttered scenes (Figure 4).
V-D Evaluation in the Simulation
Each object is grasped 20 times in the simulation experiment. We demonstrate the simulation results through the success rate and runtime in Table I.
Baselines: we compare with the following state-of-the-art baselines: 1. Heuristic grasp sampler: a heuristic grasp sampler to generate grasps based on the normal of object point clouds [12]; 2. A cVAE-based approach, FFHNet [12]; 3. A GANs-based approach, DexGanGrasp [20]; 4. A Diffusion-based approach, DexDiffuser [8].
In Table I, there is a clear performance drop from similar to novel objects, indicating the necessity of having the introspective capability to identify OOD objects. Regarding Q2, FFHFlow-lvm outperforms all baselines on both similar objects and novel objects and is ten times faster than the diffusion baseline. These results also answer Q1 that FFHFlow-lvm is able to generate high-quality and relatively fast grasp synthesis. It is attributed to its informative latent features and uncertainty-aware grasping mechanism. Notably, the diffusion-based approach [8] also demonstrates relatively better performance compared to cVAE [12] on similar objects and GAN [20] on both similar and novel objects. However, the iterative de-noising steps in diffusion models result in a significantly longer runtime ($1610$ms). In contrast, both cVAE [12] and GAN [20], owning to their architectural simplicity, are the fastest methods ($30$ms).
@@ V-D Evaluation in the Simulation
Figure 5: Analysis of shape uncertainty on object level by flow likelihoods (Top and Middle) and grasp evaluator scores (Bottom) for In-Distribution (Blue) and Out-of-Distribution Objects (Red). Flow likelihoods can better detect novel objects with low likelihoods, showcasing the introspective capability for the object shapes of the operated objects than the grasp evaluator (Bottom).
V-E Shape-Aware Introspection Analysis
V-E1 Introspection Quantification
We quantify the shape-aware introspection via the likelihoods of Prior Flow and Grasp Flow. Firstly, in Figure 5, Prior Flow is able to represent the object-level shape uncertainty. Because Prior Flow can better estimate the density of known object geometries as it only conditions on the object point clouds, this capability allows it to better distinguish between in-distribution objects and out-of-distribution (OOD) novel ones. We anticipate potential applications for the Prior Flow likelihood in scenarios with multiple objects. In these situations, the robot could prioritize grasping in-distribution objects first, as it is more confident in its ability to manipulate them. On the other hand, Grasp Flow represents the view-level shape uncertainty as it considers both the partial point cloud and the corresponding grasps (Figure 6). To showcase the effects of the introspection, we present the relationship between Grasp Stability/Grasps in Collision and different metrics, including likelihoods and the grasp evaluator scores in Figure 7.
We collect an evaluation set and generate 100 grasp candidates for each partial view. For each grasp, we obtain the likelihoods of Grasp Flow and Prior Flow, as well as the evaluator score. To assess the quality of the generated grasps, we utilize the Flexible Collision Library (FCL) to predict collisions for each grasp and Gazebo to evaluate the grasp stability of the remaining non-collided grasps. In Figure 7, we observe a clear negative correlation between the grasp evaluator score and the failure rate due to collision. In contrast, Prior Flow and Grasp Flow demonstrate the ability to reduce collision. In the bottom plot, both the evaluator score and Grasp Flow likelihoods exhibit a strong correlation with grasp stability. The grasp evaluator outperforms Grasp Flow as it was specifically trained to distinguish positive grasps from negative ones. However, the Prior Flow, representing the object-level shape uncertainty, is less relevant to grasp stability. This explains the performance drop with Prior Flow in Table II.
Figure 6: Exemplar visualization of view-level shape uncertainty (Grasp Flow likelihoods). Given the partial point cloud in red, we color the predicted grasps according to Grasp Flow likelihoods. The likelihood is lower for the grasps (in purple) closer to the unseen and hence incomplete face of the point clouds, indicating the introspective capability for the incomplete shape due to partial observation.
Figure 7: Relationship between Grasp Stability (Bottom)/ Grasps in Collision (Top) and different metrics. Grasp flow (Blue) can represent view-level shape uncertainty, striking a balance between grasp stability mainly captured by the grasp evaluator (Green) and avoiding collisions.
V-E2 Shape-Aware Introspective Grasp Evaluation
We apply Grasp Flow in our final grasping pipeline because Grasp Flow can better strike the trade-off between grasp stability and collision avoidance evidenced in Table II. The effect of Grasp Flow in filtering grasps can be illustrated based on Figure 6. Grasp candidates targeting the unseen backside of objects are assigned lower likelihoods, whereas those targeting the top and front surfaces are predicted with higher likelihoods. This demonstrates Grasp Flow’s capability to quantify shape uncertainty for various predicted grasp orientations, effectively identifying the observable regions based on likelihoods.
Figure 8: Predicted dexterous grasps in real-world experiments
TABLE II: Shape-Aware Introspective Grasp Evaluation
Fusion Strategy Similar Novel
@@ V-E2 Shape-Aware Introspective Grasp Evaluation
+ Evaluator $90.5\%$ $50.9\%$
Evaluator + Prior Flow $87.7\%$ $52.4\%$
Evaluator + Grasp Flow 94.6% 52.7%
TABLE III: Variants of Proposed Flow-based Samplers
Methods (with Evaluator only) Similar Novel Run time (ms) FFHFlow-cnf $85.4\%$ $36.7\%$ $70$ FFHFlow-lvm -light $86.7\%$ $48.5\%$ 60 FFHFlow-lvm 90.5% 50.9% $130$
V-F Evaluation in the Real World
TABLE IV: Real-World Experiment Results

### v3

@@ 4.3.1 Learning Grasp Distribution via DLVMs
Variational Lower Bound: With Jensen Inequality, we have the following variational lower bound with an approximate posterior of the latent variable $q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{g})$ (more details in the appendix):
$\begin{split}\log{p_{\theta}(\mathbf{g}|\mathbf{x})}\geq&\mathbb{E}_{q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{g})}[\log{p_{\theta}(\mathbf{g}|\mathbf{x},\mathbf{z})}]-\beta KL(q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{g})||p_{\theta}(\mathbf{z}|\mathbf{x})).\end{split}$ (3)
@@
where $\beta$ is a hyperparameter to control the extend of shape-aware information in the latents. In practice, the KLD term is implemented as the sum of the negative entropy of the approximate posterior and the cross-entropy between the prior and the approximate posterior for convenience. The effect of minimizing this term is to match the input-dependent prior $p_{\theta}(\mathbf{z}|\mathbf{x})$ to the approximate posterior $q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{g})$, and meanwhile encourage the approximate posterior to be more diverse. A flexible prior is of vital importance at inference time as we resort to the Prior Flow for drawing latent samples. These samples will be conditioned on the Grasp Flow for grasp generation (Figure 2). Therefore, an input-dependent and learnable prior is not only useful for increasing expressivity but also necessary for generating diverse grasps given an incomplete point cloud.
@@
Grasp Flow $p_{\theta}(\mathbf{g}|\mathbf{x},\mathbf{z})$: Both VAE and GANs are reported to suffer from the mode collapse problem [5], which is also named Information Preference Property [4]. It illustrates the inability of VAE to capture the entire data distribution, e.g., the complex multi-modal grasp distribution. Whenever this happens, the latent variable $\mathbf{z}$ is neglected by the powerful decoder and hence is uninformative in terms of the data, i.e., grasps and object point clouds. It has been proved that the unbounded likelihood function is the crux of this problem, e.g., an isotropic Gaussian [3]. To mitigate this, we propose to learn an cNF for the likelihood function $p_{\theta}(\mathbf{g}|\mathbf{x},\mathbf{z})$ by abuse of notations in Section 4.2: $p_{\theta}(\mathbf{g}|\mathbf{x},\mathbf{z})=p_{u|\mathbf{z}}(T_{\theta|\mathbf{z}}^{-1}(\mathbf{g};\mathbf{z});\mathbf{z})|\det(J_{T_{\theta|\mathbf{z}}^{-1}}(\mathbf{g};\mathbf{z}))|$, where the base distribution $p_{u|\mathbf{z}}:\mathcal{R}^{d}\times\mathcal{R}^{l}\rightarrow\mathcal{R}$ is conditional on $\mathbf{z}$. For conciseness, $\mathbf{x}$ is not shown on the RHS as $\mathbf{z}$ already subsumes the information from $\mathbf{x}$, so Grasp Flow is simplified as $p_{\theta}(\mathbf{g}|\mathbf{z})$. In contrast to cVAE, there is no underlying assumption for the distribution form in NFs, which allows the model to learn a more general likelihood function instead of an isotropic Gaussian. Meanwhile, the architecture of NFs is restricted to be a diffeomorphism, which we anticipate to help alleviate the unboundness issue.
@@
Prior Flow $p_{\theta}(\mathbf{z}|\mathbf{x})$: The overly-simple prior can induce excessive regularization, limiting the quality of the latent representation [6, 7], which can be observed in cVAE-based approaches [1, 12, 2] with an input-independent isotropic Gaussian as the prior (Figure 3(b)). On the other hand, the assumption of input-independence for the prior [4] poses difficulty in learning informative latent features. To address this, we propose to utilize a second cNF for an input-dependent prior distribution, in our case, a point cloud-dependent prior: $p_{\theta}(\mathbf{z}|\mathbf{x})=p_{u}(T_{\theta|\mathbf{x}}^{-1}(\mathbf{z};\mathbf{x}))|\det(J_{T_{\theta|\mathbf{x}}^{-1}}(\mathbf{z};\mathbf{x}))|$.
Variational Network $q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{g})$: The variational network is designed to approximate the real but intractable posterior distribution $p_{\theta}(\mathbf{z}|\mathbf{x},\mathbf{g})$ for amortized variational inference. $p_{\theta}(\mathbf{z}|\mathbf{x},\mathbf{g})$ is defined within the DLVM according to the Bayes formula: $p_{\theta}(\mathbf{z}|\mathbf{x},\mathbf{g})=\frac{p_{\theta}(\mathbf{g}|\mathbf{z},\mathbf{x})p_{\theta}(\mathbf{z}|\mathbf{x})}{p_{\theta}(\mathbf{g}|\mathbf{x})}$, where ${p_{\theta}(\mathbf{g}|\mathbf{x})}=\int{p_{\theta}(\mathbf{g}|\mathbf{z},\mathbf{x})p_{\theta}(\mathbf{z}|\mathbf{x})d\mathbf{z}}$ is the model evidence. From a pragmatic perspective, this network should be a powerful feature extractor for the grasps and object point clouds. As it is not used during inference, we keep it simple and use DNNs to predict a factorized Gaussian for the variational posterior distribution on the latent space $\mathbb{R}^{l}$, i.e., $q_{\phi}(\mathbf{z}|\mathbf{x},\mathbf{g})=\mathcal{N}(\mathbf{z};\mu_{\phi}(\mathbf{x},\mathbf{g}),diag(\delta_{\phi}(\mathbf{x},\mathbf{g})))$.
@@
Grasp Generation and Shape-Aware Introspection: For a test object point cloud $\mathbf{x}^{*}$, we can generate the corresponding grasps $\mathbf{g}^{*}$ by performing ancestral sampling:
@@
From Section 2 we know that we can compute the exact likelihoods of $p_{\theta}(\mathbf{g}|\mathbf{z}^{*})$ and $p_{\theta}(\mathbf{z}|\mathbf{x}^{*})$ from Grasp Flow and Prior Flow, respectively. We expect these two quantities to capture the knowledge gained by the model quantitatively. With these, our model can be introspective against its unknown knowledge, such as the incomplete shape due to partial observation and unknown object shapes due to limited coverage of training data. Specifically, the likelihoods of Grasp Flow quantify the inverse view uncertainty as it sees both the point cloud and the corresponding grasps through the learned latent variable. Thereby, high values are assigned to grasps towards visible views and low for grasps to the invisible views. Complementarily, with the proposed DLVMs, Prior Flow is expected to represent the inverse object uncertainty by purely conditioning on the input point cloud, i.e., low likelihoods for unknown object shape and high for known ones.
4.4 Uncertainty-Aware Grasp Evaluation
@@
To further secure the grasp success, we train a discriminative grasp evaluator $f_{\psi}(\mathbf{g},\mathbf{x})$ that outputs a score to better capture the grasp quality by learning to distinguish feasible grasps from infeasible ones in a supervised manner [12]. Furthermore, as Grasp Flow is able to quantify the view uncertainty, it is beneficial to incorporate this information into the grasp evaluation. With this, we can penalize the grasps approaching the uncertain side of the partially observed object to avoid potential collisions. To this end, we introduce an uncertainty-aware grasp evaluation strategy that fuses the batch-normalized view uncertainty into the grasp evaluator:
$\epsilon f_{\psi}(\mathbf{g}^{*},\mathbf{x}^{*})+(1-\epsilon)\log{p_{\theta}(\mathbf{g}^{*}|\mathbf{z}^{*})},$ (5)
where $\epsilon\in[0,1]$ is a parameter for balancing the effects of grasp quality and reducing potential collisions due to the partial observability (an ablation study in Section 3.2 of the appendix).
5 Experiment
@@
In this section, we perform comprehensive experiments and analysis to answer the following questions: Q1: Can the proposed FFHFlow-lvm facilitate more expressive latent representations for both diverse, high-quality and fast grasp synthesis? Q2: How well does FFHFlow-lvm perform compared with other state-of-the-art grasp synthesis approaches? Q3: How effective is the shape-aware introspection of FFHFlow-lvm against OOD data and does it benefit grasp evaluation? Q4: Can the model generalize to the complex real-world scenarios (e.g., confined and cluttered scenes)? More ablation studies are included in appendix due to space limit.
5.1 Experimental Setup
@@
The experiments of grasping single tabletop objects are performed in both the simulation and the real world with the DLR-HIT II hand [56]. In the real-world, except for single object grasping in unconfined space, we validate our approach in another two complex ones, i.e., the confined and cluttered scenarios (Figure 4(a)). The success criterion is defined as the ability to lift the object 20 cm above its resting position without slippage.
(a) Objects and setups evaluated in the simulation and real world. We test simulation objects with similar (1) and novel (2) shapes, real objects (3), and grasping in confined space (4) and cluttered scene (5-6).
@@
(b) Analysis of view and object uncertainty (Top and Middle) and grasp evaluator scores (Bottom) for In-Distribution (Blue) and Out-of-Distribution Objects (Red). Y axis: frequency/counts, X axis: scores/likelihoods (averaged over predicted grasps for Prior Flow and grasp evaluator). Flow likelihoods can better detect novel objects with low likelihoods.
Figure 4: Experiment Setup (a) and Object Uncertainty Evaluation (b).
Table 1: Results in Simulation (a) and Ablation Study on Uncertainty-aware Grasping (b).
Methods
Similar
Novel
Run time (ms)
Heuristic $20.9\%$ $11.1\%$ $3387$
cVAE [12] $84.6\%$ $52.4\%$ 30
GAN [21] $86.0\%$ $49.4\%$ 30
Diffusion [8] $88.2\%$ $51.7\%$ $1610$
FFHFlow-cnf $85.4\%$ $36.7\%$ $70$
FFHFlow-lvm 94.6% 52.7% $130$
Evaluation Strategy
Similar
Novel
@@ Methods
+ Evaluator $\% 50.950.9\%$
Evaluator + Prior Flow
$87.7\%$ $52.4\%$
Evaluator + Grasp Flow
94.6% 52.7%
Data: We use only simulated data generated based on a heuristic grasp planner for training in Gazebo Simulator [12]. For training, we use 77 graspable objects filtered from KIT [57] datasets based on their graspability and object type. For testing, we select objects of two levels of difficulty from the KIT and YCB dataset [58] in simulation (Figure 4(a)): (1) Similar: 12 objects from KIT dataset with similar shapes to training objects, serving as ID objects. (2) Novel: 9 objects from YCB dataset with shapes distinct from training objects, often more difficult to grasp, serving as OOD objects. For real-world evaluation, we use 12 unknown objects from YCB dataset. (More detail about the training data and implementation in Section 2 of the appendix.)
5.2 Evaluation in the Simulation
We demonstrate the simulation results through the success rate and runtime in LABEL:tab:grasping_simulation_succ. Baselines: 1. Heuristic grasp sampler: a heuristic grasp sampler to generate grasps based on the normal of object point clouds [12]; 2. A cVAE-based approach, FFHNet [12]; 3. A GANs-based approach, DexGanGrasp [21]; 4. A Diffusion-based approach, DexDiffuser [8]. We apply the same evaluator for all generative grasping baselines except for FFHFlow-lvm with the proposed uncertainty-aware evaluation strategy. Results: To answer Q1 and Q2, in LABEL:tab:grasping_simulation_succ, first there is a clear performance drop from similar to novel objects, indicating the necessity of having the introspective capability to identify OOD objects. FFHFlow-lvm is able to generate high-quality and relatively fast grasp synthesis, outperforming the baselines on both similar objects and novel objects and is 10x faster than the diffusion baseline. Notably, the diffusion-based approach [8] demonstrates relatively better performance compared to cVAE [12] on similar objects and GAN [21] on both similar and novel objects. However, the iterative de-noising process results in a significantly longer runtime (16101610ms).
5.3 Uncertainty-Aware Grasp Evaluation
(a) Visualizing view uncertainty. Predicted grasps colored based on Grasp Flow likelihoods: lower for grasps (purple) to the incomplete face of the partially perceived point cloud.
(b) Number of collided (Top) and unstable (Bottom) grasps filtered with an increasing threshold (higher the better). Likelihoods from Grasp Flow (Blue) achieve a more optimal balance between grasp stability and collision.
Figure 5: View Uncertainty Visualization (a) and Evaluation (b).
Uncertainty Quantification:
To answer Q3, Figure 4(b) shows that Prior Flow can better represent the object uncertainty via estimating the density of object shapes. It can better distinguish between in-distribution (ID) and novel/OOD objects with distinct shapes to the training data. This is useful in situations with multiple objects, where the robot could prioritize grasping in-distribution objects first to avoid grasp failures. On the other hand, Grasp Flow can represent the view uncertainty (visualized in Figure 5(a)). To showcase its effects on the predicted grasps, we present the relation between the number of collided and unstable grasps and increasing thresholds of the likelihoods and the grasp evaluator scores (higher the better) in Figure 5(b). We assess the collision of the generated grasps with the Flexible Collision Library (FCL) and the grasp stability with Gazebo simulation (more implementation details in the appendix). We observe that, in contrast to the grasp evaluator score, both Prior Flow and Grasp Flow demonstrate the ability to reduce collision in the top plot. In the bottom plot, only the evaluator score and Grasp Flow can help improve grasp stability while the Prior Flow is less effective, which also explains the performance drop in LABEL:tab:ablation_uncertainty_succ.
Uncertainty-Aware Grasp Evaluation: Grasp Flow likelihoods are utilized for grasp evaluation as it represents the view uncertainty and exhibits a better trade-off between grasp stability and collision avoidance in Figure 5(b). This capability facilitates grasp selection with higher stability and less collision, leading to greater performance gain (LABEL:tab:ablation_uncertainty_succ). The experiment in LABEL:tab:ablation_uncertainty_succ follows the same set up as LABEL:tab:grasping_simulation_succ and more details are described in Appendix Section 2.4.
5.4 Evaluation in the Real World
Table 2: Real-World Experiment Results